OpenStudy (anonymous):

Solve this initial value problem... y' = x^3 (1 - y) where y(0) = 3

OpenStudy (anonymous):

\[\int\limits\limits_{}^{}dy/(1-y) = \int\limits\limits_{}^{}x^{3}dx\]

OpenStudy (anonymous):

\[\ln 1 -y = x ^{4}/4 + C\]

OpenStudy (anonymous):

\[1-y = \exp(x ^{4}/4) + expC\]

OpenStudy (anonymous):

\[1-y = C \exp(x ^{4}/4)\]

OpenStudy (anonymous):

y(0) = 3, C=-2 \[1 + 2\exp(x ^{4}/4)=y\] But the back of the book shows -(x^4/4). Where did I go wrong??

OpenStudy (turingtest):

this is linear and I think an integrating factor will work

OpenStudy (anonymous):

At the second line the result of integral of dy ( 1-y) = - ln ( 1-y)

OpenStudy (anonymous):

yes... the dreaded minus sign

OpenStudy (anonymous):

You're right. Thank you

OpenStudy (turingtest):

\[y'=x^3(1-y)=x^3-yx^3\]\[y'-x^3y=x^3\]\[\mu=e^{\int x^3dx}=e^{x^4/4}\]haha I went the long way but it will get the same result :P

OpenStudy (anonymous):

Yeah, separate and integrate should work.

OpenStudy (turingtest):

yeah. much easier than an IF

OpenStudy (anonymous):

But, you're a champ for finding alternatives.

OpenStudy (turingtest):

thank ya kindly

OpenStudy (anonymous):

IF confuses me

OpenStudy (turingtest):

you'll get there

OpenStudy (anonymous):

I haven't seen the proof so i feel like the theorem comes out of thin air.

OpenStudy (turingtest):

mystery solved, though it's not an easy proof to memorize http://tutorial.math.lamar.edu/Classes/DE/Linear.aspx

OpenStudy (anonymous):

Thanks!

OpenStudy (turingtest):

welcome!